OpenStudy (anonymous):

What could the formula 2a+2b represent

OpenStudy (anonymous):

area of a parallelogram with side lengths a and b B. perimeter of a parallelogram side lengths a and b C. area of a trapezoid with height and base a, and side length b D. perimeter of a trapezoid with height and base a, and side length b

TheSmartOne (thesmartone):

Do you know what the area and perimeter of a rectangle is?

OpenStudy (anonymous):

what if you mean the formulas for them then ye s

TheSmartOne (thesmartone):

Write the formulas...

OpenStudy (anonymous):

C

OpenStudy (anonymous):

A of a rectangle = ab and P= 2* l+w

TheSmartOne (thesmartone):

Perimeter= 2(l*w)=2l+2w

TheSmartOne (thesmartone):

which means b not C. -_-

TheSmartOne (thesmartone):

@imasim

OpenStudy (anonymous):

thanks bro